OpenStudy (anonymous):

Find x^2 + y^2+ z^2 - 4xyz if \[\sin ^{-1}x+\sin ^{-1} y+\sin ^{-1}z = \frac{ 3\Pi }{ 2 }\]?

OpenStudy (shubhamsrg):

hint : max value of sin^-1 x ?

OpenStudy (anonymous):

ye sawal kaha se liya hai

OpenStudy (shubhamsrg):

arre,,arent you getting the catch ? that was a great thing acc to me,,whats the max value of xin^-1 x? o.O

OpenStudy (shubhamsrg):

great hint* :P

OpenStudy (anonymous):

u mean max value of sinx-1 or sinx

OpenStudy (shubhamsrg):

sin^-1

OpenStudy (shubhamsrg):

inverse

OpenStudy (anonymous):

pi/2 + pi/2 +pi/2 do u mean that

OpenStudy (shubhamsrg):

yep..the only possible ans..

OpenStudy (anonymous):

hmm i don't think so

OpenStudy (shubhamsrg):

really,,what makes you argue ?

OpenStudy (anonymous):

try that formula sin-1x + sin-1y=sin-1(x(1-y^2)^1/2+y(1-x^2)^1/2)

OpenStudy (anonymous):

bring sins into one form like sin-1(.....)=3pi/2

OpenStudy (anonymous):

ok u are may be right

OpenStudy (anonymous):

i was just trying to confirm it

OpenStudy (shubhamsrg):

ofcorse am right,, suppose i ask you sinx +siny + sinz =3, then ofcorse its possible when all are =1

OpenStudy (shubhamsrg):

same with this case.. hence x=y=z =1

OpenStudy (anonymous):

@shubhamsrg Why Did u Take sin as Maximum ? Sorry My Connection was Out Of Order :)

OpenStudy (anonymous):

@satellite73 @RadEn

OpenStudy (anonymous):

@shubhamsrg provided you with the answer

OpenStudy (anonymous):

i go with @shubhamsrg\[\sin ^{-1}x+\sin ^{-1} y+\sin ^{-1}z \le \frac{ 3\pi }{ 2 }\]equality occurs when\[\sin ^{-1}x=\sin ^{-1} y=\sin ^{-1}z = \frac{\pi }{ 2 }\]

OpenStudy (anonymous):

you have \[\sin^{-1}(x)+\sin^{-1}(y)+\sin^{-1}(z)=\frac{3\pi}{2}\] but the the very largest \(\sin^{-1}(x)\) can be is \(\frac{\pi}{2}\) so they must each be \(\frac{\pi}{2}\), otherwise they cannot add to \(\frac{3\pi}{2}\)